(A) two unbiased dice are thrown.
find the probability that they
(i) show the same number
(ii) show different number
(B) three unbiased dice are thrown.
find the probability that
(i) they all show different
numbers
(ii) at least two show the same
number
please kindly explain how u guys get the ans. 
i suck at probability
(ai) 1/6,
do tree diagrams or whatever you want to do to find the
scenarios like
(1,1) , (2, 2), (3, 3) and so on.. each scenario probability =
1/36,
there's 6 of these scenarios so 6 x 1/36 = 1/6
(btw it's 1/36 + 1/36 + 1/36 ...)
(aii) ans = 1 - ans in (ai) = 1 - 1/6 = 5/6